OpenStudy (anonymous):

At 25 degree celsius, the equilibrium constant Kc for the reaction 2A(g) yields B(g) + C(g) is 0.035. A mixture of 8.00 moles of B and 12.00 moles of C in a 20.0 L container is allowed to come to equilibrium. What is the equilibrium concentration of A? A. < 0.100 B. 13.56 C. 0.678 D. 0.339 E. 6.78

OpenStudy (aaronq):

write the expression, plug in your values, and solve.

OpenStudy (anonymous):

I did but the answer I got wasn't what is given

OpenStudy (aaronq):

i forgot to mention that you need an ice table

OpenStudy (anonymous):

This is what I have for my ICE table. 2A(g) yields B(g) + C(g) Initial 0 0.400 0.600 Change -2x +x +x Equili -2x 0.400+x 0.600+x Kc = (B)(C)/(A)^2 0.035 = 0.400(0.600)/x^2 0.035(x^2) = 0.240 x^2 = 6.857 x = 2.618587

OpenStudy (aaronq):

you need to reverse the signs, you're losing B and C and gaining A 2A(g) yields B(g) + C(g) Initial 0 0.400 0.600 Change +2x -x -x Equili +2x 0.400-x 0.600-x

OpenStudy (anonymous):

Oh. I see my mistake

OpenStudy (anonymous):

0.035(2x) = 0.400(0.600) 0.070x = 0.240 x = 3.42857 Is the answer I got but its still not the answer. Awwwwww!

OpenStudy (anonymous):

I think I figure it out. Just plug in the answer back into the Equilibrium part of the ICE table. So 2(3.43) = 6.84

OpenStudy (aaronq):

it should be: \(0.035=\dfrac{[0.400-x][0.600-x]}{[2x]^2} \)

OpenStudy (aaronq):

{x=0.3385975230485919, x=0.8241931746258267} 2(0.82)=1.64 2(0.33)=0.66